# Freeze-in and Freeze-out in Right-Handed Neutrino Extended MSSM with Seesaw

### v1

@@ Abstract
We investigate the possibility of saturating the relic density bound with light higgsinos. When the minimal supersymmetric Standard Model is extended with right-handed neutrino superfields and the seesaw scale is very low, right-handed sneutrinos can be produced via the freeze-in mechanism. In such a case we can have essentially two independent sources for dark matter, the traditional freeze-out of higgsinos and the freeze-in of right-handed sneutrinos. The heavier of these two will decay to the lighter species with a delay. We rule out such a scenario for all seesaw models as the lifetime of sterile neutrinos, produced over-abundantly via Dodelson–Widrow mechanism, exceeds the age of the universe and will contribute to the relic density.
@@ I Introduction
Supersymmetry with R-parity conservation always provides a dark matter candidate, namely the lightest supersymmetric particle (LSP). The most studied candidate is the lightest neutralino, which would act as a weakly interacting massive particle (WIMP). However, recent direct detection experiments [2, 3, 4] have excluded nearly all of the parameter space for WIMP-like dark matter candidates. In addition, both higgsino and wino lead to an underabundance unless their mass is about $1$ TeV for higgsinos and $2$ TeV for winos [5]. This leads to some level of fine-tuning in the electroweak symmetry-breaking condition of the MSSM:
@@
where $m_{H_{u/d}}^{2}$ are soft supersymmetry breaking masses and $|\mu|$ is essentially the higgsino mass. A priori, the supersymmetry breaking parameters and the $\mu$-parameter are independent. They should therefore not be too far above the electroweak scale, otherwise large cancellations are needed between them. Hence, natural supersymmetry prefers a light higgsino, while other superpartners can be heavier [6, 7]. This motivates the exploration of alternatives to WIMP dark matter.
An alternative way to produce dark matter is the freeze-in mechanism, where the dark matter candidate is so weakly coupled to other particles that it never attains thermal equilibrium [8, 9]. This requires the couplings between dark matter and the thermal bath to be of the order $10^{-7}$ or below. Such small couplings arise naturally in electroweak scale seesaw models if the MSSM is extended with sterile neutrino superfields. This could then create a possible link between neutrino physics and dark matter. For example, if the lightest sneutrino is lighter than the lightest superpartner in the thermal bath, there will always also be a freeze-out contribution, since the frozen-out particle would eventually decay to the lightest sterile sneutrino state. A sterile sneutrino with tiny Yukawa couplings would evade all direct detection constraints.
Such a scenario could be compatible with light higgsinos, as suggested by the absence of fine-tuning in Eq. 1. In this framework, while the sterile sneutrino serves as the dark matter candidate, one could also consider the higgsino playing that role with the freeze-in contribution from sterile sneutrinos giving an extra contribution to yield the correct amount of dark matter. However, this scenario is now practically ruled out by current direct detection constraints [2, 3, 4] on the spin-independent DM-nucleon scattering cross-section.
On the model side, there are four main options for adding the right-handed neutrino superfields. First, one can add them with the corresponding Yukawa couplings, leading to Dirac neutrinos. If we include the Majorana mass term for the right-handed neutrino, we get the type-I seesaw mechanism [10, 11]. The other two possibilities involve two sets of sterile neutrinos, giving the linear [12] and inverse seesaw [13] models. In both cases, we have a small lepton number violating term, which makes the active neutrino masses small. By contrast, seesaw mechanisms of type-II or type-III cannot produce freeze-in DM candidates as the (s)neutrinos have gauge interactions that lead to their thermalisation.
@@
In this work, we study models where right-handed neutrinos couple feebly to the MSSM and the lightest higgsino mass lies at the electroweak scale, as suggested by naturalness arguments. We find that the case of higgsino DM is severely constrained by direct detection experiments. We show that the freeze-in scenario requires very light sterile neutrinos. These neutrinos are so long-lived that they themselves become a component of dark matter. However, in the region of the parameter space where the freeze-in mechanism does not lead to overabundance, sterile neutrino production via the Dodelson–Widrow mechanism [19] will lead to overabundance in all simple SUSY seesaw models. Therefore, only the Dirac neutrino model, where sterile neutrinos mass eignestates are absent, can survive the relic density constraints [18].
Our paper is organized as follows. In Sec II, we review the MSSM extended with various seesaw mechanisms and constraints on them, Sec III discusses the freeze-out and freeze-in processes generating dark matter and the properties of the dark matter candidates. In Sec IV, we outline our numerical procedure and present the results and then conclude in the following section.
@@ II Seesaw extensions of the MSSM
The freeze-in mechanism requires the DM candidate to remain out of thermal equilibrium. This implies that all its couplings to SM particles must be tiny, below $10^{-7}$ or so. Since, all gauge interactions woull lead to rapid thermalisation, models in which (s)neutrinos are in a nontrivial representation of the SM gauge group cannot provide a freeze-in DM candidate. For this reason, we do not consider type-II and type-III seesaw mechanisms, where the MSSM field content is extended by SU(2) triplet superfields.
@@ II.1 Models with sterile (s)neutrinos
The simplest of models with sterile neutrino superfields is that of Dirac neutrinos. We extend the MSSM with three right-handed neutrino superfields $N^{c}_{i}$ (with lepton number L=1L=-1), which are singlets under the SM gauge group. We write the superpotential in the form
@@
Adding a Majorana mass term to this superpotential,
$W=W_{\mathrm{Dirac}}+\frac{1}{2}M_{Ni}N_{i}^{c}N_{i}^{c},$ (4)
@@
which results in small neutrino Yukawa couplings $y^{\nu}\lesssim 10^{-10}$. The extreme case $M_{N}\rightarrow 0$ reduces to the Dirac neutrino scenario. Since the right-handed (s)neutrinos are coupled to other particles only via the neutrino Yukawa couplings, they decouple from the thermal bath in the early Universe and they become rather long-lived or stable. Hence, their abundance is determined through the freeze-in mechanism.
Linear and inverse seesaw models are obtained by adding a second set of sterile neutrino superfields $S_{i}$ (with lepton number $L=1$), resulting in the superpotentials
@@
The approximate formulae for light neutrino masses are:
@@
$m_{\nu}$ $\simeq$ $\left(\frac{y^{\nu}v\sin\beta}{\sqrt{2}M_{N}}\right)^{2}\mu_{S},$ (10)
for linear and inverse seesaws, respectively. Compared to type-I seesaw, these differ by factors of $y^{LS}/y^{\nu}$ and $\mu_{S}/M_{N}$ multiplying the mass formula.
As we shall discuss later, in the region of parameter space that corresponds to an acceptable relic density, the sterile neutrinos in the type-I seesaw are so light that they become cosmologically stable and constitute an additional dark matter component. The relic abundance generated via the Dodelson–Widrow mechanism [19] depends on the mixing angle between active and sterile neutrinos, and for type-I seesaw, we will show that sterile neutrino production leads to overabundance. Therefore, we also study linear and inverse seesaw mechanisms, as they could suppress this mixing as the relation between active and sterile neutrino masses is different.
@@ II.2 Sneutrinos and their interactions
In the presence of a seesaw mechanism, the mass matrix of the sneutrinos contains a lepton number violating part leading to a mass splitting between the CP-even and CP-odd sneutrino components [20, 21]. In general, the sneutrinos are produced in states of a definite lepton number. If the mass splitting between the sneutrino components is small, both CP-components propagate coherently and the state oscillates between a sneutrino state and an antisneutrino state [20, 21, 22], with the oscillation period being proportional to $\delta m^{2}$, the mass splitting between the CP-even and CP-odd sneutrino states.
@@
The terms of the form $A_{\nu}v_{u}-y^{\nu}\mu v_{d}$, which lie outside the $2\times 2$ block-diagonal parts of the mass matrix, mix left- and right-handed sneutrinos. In principle, this mixing could be large due to large A-terms [23], but in that case right-handed sneutrinos would thermalise through the gauge interactions of the left-handed component. For simplicity, we assume that $A_{\nu}$ to be vanishingly small, which also makes the mixing between left- and right-handed neutrinos negligible.
From the superpotential of Eqs. 4, 7 and 8, we see that the right-handed sneutrinos only interact with Higgs bosons, higgsinos, leptons and sleptons. The vertices between $\tilde{N}$ and $\tilde{H}^{\pm}\ell^{\mp}$, $\tilde{H}^{0}\nu$ are given by $y^{\nu}$, while in all but linear seesaw the vertices between the Higgs bosons and sneutrinos are
@@
and the ones for charged Higgs, charged sleptons and right-handed sneutrinos are those of Eqs. 13 and 15 with an opposite sign. Since $y^{\nu}$ is extremely small and we assume $M\ll\mu$, we may neglect the interactions of Eqs. 12 and 14.
@@
As alluded to above, the active-sterile mixing in this scenario becomes problematic because of the long lifetime of the sterile neutrino. In principle, one could increase $y^{LS}$ to achieve a viable neutrino spectrum with a heavier sterile neutrino in order to suppress the active-sterile mixing to an acceptable level for sterile neutrino production. However, this would imply a larger coupling between the Higgs boson and sneutrinos, which increase the freeze-in contribution of sneutrinos and leads to overabundance. Hence, in the linear seesaw, it is impossible to suppress the active-sterile mixing in the neutrino spectrum, keep a viable neutrino mass spectrum and satisfy the relic density constraints simultaneously.
III Relic density and direct detection properties
@@ III.1 Freeze-out component
In either case, the heavier of the two will eventually decay into the lighter species. However, due to the feeble interactions between the two sectors, the decay occurs with a delay. Typically, this delay is so long that by the time it happens, the number density of the lighter species has already diluted so much that the decay products no longer result in any further annihilations. As a result, the freeze-out and freeze-in processes remain independent.
@@
The higgsino multiplet consists of two neutral higgsinos and one charged higgsino, all with masses close to $|\mu|$. All of them mix with gauginos, although the mixing is small if $|M_{1}|,|M_{2}|\gg|\mu|$. The annihilation of higgsinos is mainly a coannihilation process of the neutral and charged higgsino through the $W$-boson, shown in Fig. 1. Since the coupling is the weak gauge coupling, the annihilation is always efficient.
@@
Figure 1: Higgsinos annihilate mainly through the $W$-portal as the coupling is unsuppressed. Here SM denotes Standard Model fermions.
A subleading contribution in the annihilation comes from the $Z$-portal. The coupling between the lightest higgsinos and the $Z$-boson is
@@
Here $N_{11}$ and $N_{12}$ denote the up- and down-type higgsino components of the lightest neutralino. When $|M_{1}|,|M_{2}|\gg|\mu|$, the mass matrix of Eq. 18 leads to $|N_{11}|$ and $|N_{12}|$ being nearly equal ($\sim 1/\sqrt{2}$) and hence, in the case of an almost pure higgsino, the coupling of Eq. 19 is strongly suppressed. As a result, the higgsino produces a much weaker signal in direct detection than one would expect from its annihilation rate.
However, even before the latest results on the spin-independent WIMP-nucleon cross section [24, 3, 4], a large part of the parameter space, even with the suppressed couplings, was already excluded. The latest LZ bound [2] on the direct detection cross section now requires the gauginos to be around $15$ TeV in order of the higgsino couplings to the $Z$ boson to be sufficiently suppressed. Such a spectrum cannot be considered natural even though it is technically allowed [^1] The heavier higgsinos would become long-lived so further studies would be needed to evaluate, to what extent the LHC searches of displaced vertices or disappearing tracks would have excluded such a scenario already. For the unphysical model of a pure higgsino the bound based on disappearing tracks is 210 GeV [25].. These constraints are obviously evaded if the sneutrino is the dark matter candidate.
@@ III.2 Freeze-in component
Right-handed sneutrinos can be produced through their interactions with both the Higgs sector and the higgsino sector. As discussed in Sec II, these interactions are proportional to the neutrino Yukawa couplings $y^{\nu}$ or the trilinear scalar couplings $A_{\nu}$. In what follows, we assume the latter to vanish.
Since Higgs bosons and higgsinos couple most strongly to the third generation of fermions, as well as to electroweak gauge bosons, and gauginos, the dominant freeze-in production arises from a limited set of $2\rightarrow 2$ processes as shown in Fig. 2. Depending on the process, the mediator can appear in either the $t$-channel or the $s$-channel, although some of these processes are kinematically suppressed.
@@
    Higgsino-sector mediators: Neutral and charged higgsinos can also mediate freeze-in production. Neutral higgsino exchange couples left-handed neutrinos and right-handed sneutrinos to third generation quark-squark or tau-stau pair (in Fig. 2(c)), while charged higgsino exchange connects charged leptons and right-handed sneutrinos to third-generation quark-squark, a tau-sneutrino or a stau-neutrino pairs (in Fig. 2(d)).
@@
Figure 2: Potential t- and s-channel processes contributing to the freeze-in production of the right-handed sneutrinos. The mediator can be (a) the SM Higgs, (b) the charged Higgs, (c) a neutral higgsino, (d) the charged higgsino, (e) a left-handed neutrino, or (f) a left-handed sneutrino. SM stands for Standard Model particles ($t$, $b$, $\tau$, $W$, $Z$).

### v2

# Freeze-in and Freeze-out in a Right-Handed Neutrino Extended MSSM with a Seesaw Mechanism

@@ Abstract
We investigate the possibility of saturating the relic density bound with light Higgsinos. When the minimal supersymmetric Standard Model is extended with right-handed neutrino superfields and the seesaw scale is very low, right-handed sneutrinos can be produced via the freeze-in mechanism. In such a case we can have essentially two independent sources for dark matter, the traditional freeze-out of Higgsinos and the freeze-in of right-handed sneutrinos. The heavier of these two will decay to the lighter species with a delay. We rule out such a scenario for all seesaw models as the lifetime of sterile neutrinos produced over-abundantly via Dodelson–Widrow mechanism exceeds the age of the universe and will contribute to the relic density.
@@ I Introduction
Supersymmetry with R-parity conservation always provides a dark matter candidate, namely, the lightest supersymmetric particle (LSP). The most studied candidate is the lightest neutralino, which would act as a weakly interacting massive particle (WIMP). However, recent direct-detection experiments [2, 3, 4] have excluded nearly all of the parameter space for WIMP-like dark matter candidates. In addition, both the Higgsino and the wino lead to an underabundance unless their mass is about $1$ TeV for Higgsinos and $2$ TeV for winos [5]. This leads to some level of fine-tuning in the electroweak symmetry-breaking condition of the minimal supersymmetric Standard Model (MSSM):
@@
where $m_{H_{u/d}}^{2}$ are soft supersymmetry-breaking masses and $|\mu|$ is essentially the Higgsino mass. A priori, the supersymmetry-breaking parameters and the $\mu$ parameter are independent. They should therefore not be too far above the electroweak scale, otherwise large cancellations are needed between them. Hence, natural supersymmetry prefers a light Higgsino, while other superpartners can be heavier [6, 7]. This motivates the exploration of alternatives to WIMP dark matter.
An alternative way to produce dark matter is the freeze-in mechanism, where the dark matter candidate is so weakly coupled to other particles that it never attains thermal equilibrium [8, 9]. This requires the couplings between dark matter and the thermal bath to be of the order $10^{-7}$ or below. Such small couplings arise naturally in electroweak-scale seesaw models if the MSSM is extended with sterile-neutrino superfields. This could then create a possible link between neutrino physics and dark matter. For example, if the lightest sneutrino is lighter than the lightest superpartner in the thermal bath, there will always also be a freeze-out contribution, since the frozen-out particle would eventually decay to the lightest sterile-sneutrino state. A sterile sneutrino with tiny Yukawa couplings would evade all direct-detection constraints.
Such a scenario could be compatible with light Higgsinos, as suggested by the absence of fine-tuning in Eq. 1. In this framework, while the sterile sneutrino serves as the dark matter candidate, one could also consider the Higgsino playing that role with the freeze-in contribution from sterile sneutrinos giving an extra contribution to yield the correct amount of dark matter. However, this scenario is now practically ruled out by current direct-detection constraints [2, 3, 4] on the spin-independent DM nucleon scattering cross-section.
On the model side, there are four main options for adding the right-handed neutrino superfields. First, one can add them with the corresponding Yukawa couplings, leading to Dirac neutrinos. If we include the Majorana mass term for the right-handed neutrino, we get the type-I seesaw mechanism [10, 11]. The other two possibilities involve two sets of sterile neutrinos, giving the linear [12] and inverse seesaw [13] models. In both cases, we have a small lepton number violating term, which makes the active neutrino masses small. By contrast, the seesaw mechanisms of type-II or type-III cannot produce freeze-in DM candidates as the (s)neutrinos have gauge interactions that lead to their thermalisation.
@@
In this work, we study models where right-handed neutrinos couple feebly to the MSSM and the lightest Higgsino mass lies at the electroweak scale, as suggested by naturalness arguments. We find that the case of Higgsino DM is severely constrained by direct-detection experiments. We show that the freeze-in scenario requires very light sterile neutrinos. These neutrinos are so long-lived that they themselves become a component of dark matter. However, in the region of the parameter space where the freeze-in mechanism does not lead to overabundance, sterile neutrino production via the Dodelson–Widrow mechanism [19] will lead to overabundance in all simple SUSY seesaw models. Therefore, only the Dirac neutrino model, where sterile neutrinos mass eignestates are absent, can survive the relic density constraints [18]. In [20] the authors studied a similar scenario than we do, but their scan resulted in a long-lived NLSP, which would affect Big Bang nucleosynthesis.
Our paper is organized as follows: In Section II, we review the MSSM extended with various seesaw mechanisms and constraints on them; Section III discusses the freeze-out and freeze-in processes generating dark matter and the properties of the dark matter candidates. In Section IV, we outline our numerical procedure and present the results and then conclude in the following section.
@@ II Seesaw extensions of the MSSM
The freeze-in mechanism requires the DM candidate to remain out of thermal equilibrium. This implies that all its couplings to SM particles must be tiny, below $10^{-7}$ or so. Since, all gauge interactions would lead to rapid thermalisation, models in which (s)neutrinos are in a nontrivial representation of the SM gauge group cannot provide a freeze-in DM candidate. For this reason, we do not consider type-II and type-III seesaw mechanisms, where the MSSM field content is extended by SU(2) triplet superfields.
@@ II.1 Models with sterile (s)neutrinos
The simplest of models with sterile-neutrino superfields is that of Dirac neutrinos. We extend the MSSM with three right-handed neutrino superfields $N^{c}_{i}$ (with lepton number L=1L=-1), which are singlets under the SM gauge group. We write the superpotential in the form
@@
Adding a Majorana mass term to this superpotential
$W=W_{\mathrm{Dirac}}+\frac{1}{2}M_{Ni}N_{i}^{c}N_{i}^{c}$ (4)
@@
which results in small neutrino Yukawa couplings $y^{\nu}\lesssim 10^{-10}$. The extreme case $M_{N}\rightarrow 0$ reduces to the Dirac neutrino scenario. Since the right-handed (s)neutrinos are coupled to other particles only via the neutrino Yukawa couplings, they decouple from the thermal bath in the early Universe, and they become rather long-lived or stable. Hence, their abundance is determined through the freeze-in mechanism.
Linear and inverse seesaw models are obtained by adding a second set of sterile-neutrino superfields $S_{i}$ (with lepton number $L=1$), resulting in the superpotentials
@@
The approximate formulae for light neutrino masses are
@@
$m_{\nu}$ $\simeq$ $\left(\frac{y^{\nu}v\sin\beta}{\sqrt{2}M_{N}}\right)^{2}\mu_{S}$ (10)
for linear and inverse seesaw models, respectively. Compared to type-I seesaw model, these differ by factors of $y^{LS}/y^{\nu}$ and $\mu_{S}/M_{N}$ multiplying the mass formula.
As we shall discuss later, in the region of parameter space that corresponds to an acceptable relic density, the sterile neutrinos in the type-I seesaw model are so light that they become cosmologically stable and constitute an additional dark matter component. The relic abundance generated via the Dodelson–Widrow mechanism [19] depends on the mixing angle between active and sterile neutrinos, and for the type-I seesaw model, we will show that sterile-neutrino production leads to overabundance. Therefore, we also study linear and inverse seesaw models, as they could suppress this mixing as the relation between active and sterile-neutrino masses is different.
@@ II.2 Sneutrinos and their interactions
In the presence of a seesaw mechanism, the mass matrix of the sneutrinos contains a lepton number violating part leading to a mass splitting between the CP-even and CP-odd sneutrino components [21, 22]. In general, the sneutrinos are produced in states of a definite lepton number. If the mass splitting between the sneutrino components is small, both CP-components propagate coherently, and the state oscillates between a sneutrino state and an antisneutrino state [21, 22, 23], with the oscillation period being proportional to $\delta m^{2}$, the mass splitting between the CP-even and CP-odd sneutrino states.
@@
The terms of the form $A_{\nu}v_{u}-y^{\nu}\mu v_{d}$, which lie outside the $2\times 2$ block-diagonal parts of the mass matrix, mix left- and right-handed sneutrinos. In principle, this mixing could be large due to large A-terms [24], but in that case right-handed sneutrinos would thermalise through the gauge interactions of the left-handed component. For simplicity, we assume that $A_{\nu}$ is vanishingly small, which also makes the mixing between left- and right-handed neutrinos negligible.
From the superpotential of Eqs. 4, 7 and 8, we see that the right-handed sneutrinos only interact with Higgs bosons, Higgsinos, leptons and sleptons. The vertices between $\tilde{N}$ and $\tilde{H}^{\pm}\ell^{\mp}$, $\tilde{H}^{0}\nu$ are given by $y^{\nu}$, while in all but the linear seesaw model the vertices between the Higgs bosons and sneutrinos are
@@
and the ones for charged Higgs bosons, charged sleptons, and right-handed sneutrinos are those of Eqs. 13 and 15 with an opposite sign. Since $y^{\nu}$ is extremely small and we assume $M\ll\mu$, we may neglect the interactions of Eqs. 12 and 14.
@@
As alluded to above, the active-sterile mixing in this scenario becomes problematic because of the long lifetime of the sterile neutrino. In principle, one could increase $y^{LS}$ to achieve a viable neutrino spectrum with a heavier sterile neutrino in order to suppress the active-sterile mixing to an acceptable level for sterile-neutrino production. However, this would imply a larger coupling between the Higgs boson and sneutrinos, which increase the freeze-in contribution of sneutrinos and leads to overabundance. Hence, in the linear seesaw model, it is impossible to suppress the active-sterile mixing in the neutrino spectrum, keep a viable neutrino mass spectrum and satisfy the relic density constraints simultaneously.
III Relic density and direct-detection properties
@@ III.1 Freeze-out component
In either case, the heavier of the two will eventually decay into the lighter species. However, due to the feeble interactions between the two sectors, the decay occurs with a delay. Typically, this delay is so long that by the time it happens, the number density of the lighter species has already diluted so much that the decay products no longer result in any further annihilations. As a result, the freeze-out and freeze-in processes remain independent. For our parameter choices (see Table 1 below) the lifetime of the NLSP becomes of the order of $0.12$ seconds, so the energetic particles released in the decay of the NLSP have time to thermalise before Big Bang nucleosynthesis.
@@
The Higgsino multiplet consists of two neutral Higgsinos and one charged Higgsino, all with masses close to $|\mu|$. All of them mix with gauginos, although the mixing is small if $|M_{1}|,|M_{2}|\gg|\mu|$. The annihilation of Higgsinos is mainly a coannihilation process of the neutral and charged Higgsino through the $W$ boson, shown in Fig. 1. Since the coupling is the weak gauge coupling, the annihilation is always efficient.
@@
Figure 1: Higgsinos annihilate mainly through the $W$ portal as the coupling is unsuppressed. Here SM denotes Standard Model fermions.
A subleading contribution in the annihilation comes from the $Z$ portal. The coupling between the lightest Higgsinos and the $Z$ boson is
@@
Here $N_{11}$ and $N_{12}$ denote the up- and down-type Higgsino components of the lightest neutralino. When $|M_{1}|,|M_{2}|\gg|\mu|$, the mass matrix of Eq. 18 leads to $|N_{11}|$ and $|N_{12}|$ being nearly equal ($\sim 1/\sqrt{2}$) and hence, in the case of an almost pure Higgsino, the coupling of Eq. 19 is strongly suppressed. As a result, the Higgsino produces a much weaker signal in direct-detection than one would expect from its annihilation rate.
However, even before the latest results on the spin-independent WIMP-nucleon cross section [25, 3, 4], a large part of the parameter space, even with the suppressed couplings, was already excluded. The latest LZ Collaboration bound [2] on the direct-detection cross section now requires the gauginos to be around $15$ TeV in order for the Higgsino couplings to the $Z$ boson to be sufficiently suppressed. Such a spectrum cannot be considered natural even though it is technically allowed. [^1] The heavier Higgsinos would become long-lived, so further studies would be needed to evaluate to what extent the LHC searches of displaced vertices or disappearing tracks would have excluded such a scenario already. For the unphysical model of a pure Higgsino, the bound based on disappearing tracks is 210 GeV [26]. These constraints are obviously evaded if the sneutrino is the dark matter candidate.
@@ III.2 Freeze-in component
Right-handed sneutrinos can be produced through their interactions with both the Higgs sector and the Higgsino sector. As discussed in Section II, these interactions are proportional to the neutrino Yukawa couplings $y^{\nu}$ or the trilinear scalar couplings $A_{\nu}$. In what follows, we assume the latter to vanish.
Since Higgs bosons and Higgsinos couple most strongly to the third generation of fermions, as well as to electroweak gauge bosons, and gauginos, the dominant freeze-in production arises from a limited set of $2\rightarrow 2$ processes as shown in Fig. 2. Depending on the process, the mediator can appear in either the $t$ channel or the $s$ channel, although some of these processes are kinematically suppressed.
@@
    Higgsino-sector mediators: Neutral and charged Higgsinos can also mediate freeze-in production. Neutral Higgsino exchange couples left-handed neutrinos and right-handed sneutrinos to third-generation quark-squark or tau-stau pair (in Fig. 2(c)), while charged Higgsino exchange connects charged leptons and right-handed sneutrinos to third-generation quark-squark, tau-sneutrino or stau-neutrino pairs (in Fig. 2(d)).
@@
Figure 2: Potential t- and s-channel processes contributing to the freeze-in production of the right-handed sneutrinos. The mediator can be (a) the SM Higgs boson, (b) the charged Higgs boson, (c) a neutral Higgsino, (d) the charged Higgsino, (e) a left-handed neutrino, or (f) a left-handed sneutrino. SM stands for Standard Model particles ($t$, $b$, $\tau$, $W$, $Z$).